Consider a makeup mirror that produces a magnification of 1.6 when a person’s face is 14 cm away.

What is the focal length of the makeup mirror in meters?
Answer #1

Given, u=-14cm. m = lob we know, m= 31 = 0 = -mu = -106x (-14) - 22.4 cm. For mirror, f = 1/3 + 4 - = N ㅓ다 22.4 14 14 2264 22Hope it helps.

Kindly like the answer and comment your doubts in the comment section below .